# Spin-Orbital Intertwined Topological Superconductivity in a Class of Correlated Noncentrosymmetric Materials

### v1

@@ Abstract
Phase diagram and sign-change SC states.— Once superconductivity is induced by antiferromagnetic fluctuations, the pairing symmetry of the superconducting state can be classified according to the IRs of the $C_{4v}$ group. Table 1 lists the possible pairings with their form factors up to the next-nearest neighbor and their corresponding IRs. By solving the linear gap equation in Eq. (10), the leading superconducting pairing can be determined.
@@
Figure 2: The phase diagrams of superconducting states as functions of some parameters. (a)-(b) $\lambda_{R}$-$\mu$ phase diagram. $U=4.5$ and $J=0.03U$ in (a). $U=5.4$ and $J=0.21U$ in (b). (c)-(d) $U$-$J$ phase diagram. $\lambda_{R}=0.9$ and $\mu=-5.3$ in (c). $\lambda_{R}=0.8$ and $\mu=-4.5$ in (d). In (a)-(d), all other parameters are the same as those in Fig. 1 (b). In panel (a), the positions marked by numbers 1 and 2 correspond to the star-marked positions labeled 1 and 2 in (c) and (d). The notation $S_{\pm}$ and $d_{\pm}$ in $A_{1}$ and $B_{2}$ indicates that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction.
The phase diagrams of the superconducting state with respect to $\lambda_{R}$-$\mu$ and $U$-$J$ are shown in Figs. 2 and Figs. S4 and S6[43]. Over a wide-range parameter space, we identify three possible leading pairing channels: $A_{1}$, $B_{1}$, and $B_{2}$. Notably, the superconducting states in the $A_{1}$ and $B_{2}$ channels exhibit highly unconventional sign-changing gap structures $A_{1}(S_{\pm})$ and $B_{2}(d_{\pm})$, which indicate that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction as shown in Fig. 3 (a) and Fig. S9 (a)[43].
Figure 3: (a)-(d) The superconducting gap functions projected onto the corresponding Fermi surfaces. The red and blue colors indicate the opposite signs of superconducting gap functions. (a) and (b) correspond to full and $A_{1,1}=i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ channels in $A_{1}(S_{\pm})$ IR, respectively. (c) The renormalized gap function $\Delta_{A_{1}}(k_{F},\theta)/max(\Delta_{A_{1}}(k_{F},\theta))$ and $\Delta_{A_{1,1}}(k_{F},\theta)/max(\Delta_{A_{1,1}}(k_{F},\theta))$. (d) The weights of six sub-channels in $A_{1}(S_{\pm})$ IR as function as $t_{2}$. (e)-(f) The fitting gap function $\Delta_{1/2}(k)$ in Eq. 13 along different high-symmetry lines $\Gamma-M$ and $\Gamma-X$. Other parameters in (a)-(d) are the same as those in Fig. 1 (c) and of star labeled 2 in Fig. 2 (a) and (d).
To elucidate the origin of this unconventional sign-change superconducting state, we take the $A_{1}(S_{\pm})$ channel as an example. The gap function of noncentrosymmetric systems can be parameterized as follows[44],
$\hat{\Delta}_{k}=[\hat{\phi}_{k}+\bm{\hat{d}_{k}}\cdot\bm{\sigma}]i\sigma^{2}.$ (12)
In the one-orbital Rashba model, the projection of the gap function onto the two spin-split Fermi surfaces can be expressed in a simplified form, $\Delta_{1/2}(k)=\phi_{k}\pm|\hat{d}_{k}|$ with $\phi_{k}$ as a scalar. Thus, the aforementioned sign-change gap structure only occur when the triplet component is relatively strong and has helical $p$- or $f$-wave spin-triplet pairing, such as when the filling is near the van Hove singularity[32, 34, 35]. Any doping away from the van Hove singularity consistently results in predominant d-wave spin-singlet pairing[33]. However, for the two-orbital model considered here, additional possibilities arise. We project the six sub-channels $i\sigma^{2}$($\tau^{0}\sigma^{0}\phi_{1k}$, $\tau^{3}\sigma^{0}\phi_{2k}$, $\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}$, $\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$, $\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}$, $\tau^{2}\sigma^{3}d_{6k}$) in the spin-orbital representation, listed in Table 1 onto the two Fermi surfaces. Note that the first two sub-channels belong to spin-singlet while the latter four correspond to spin-triplet. The projection of the first sub-channel $i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ is shown in Fig. 3 (b), while the others are presented in Fig. S7 [43]. We find that their superconducting gap functions exhibit the same projected distribution on the Fermi surfaces, ensuring consistency with the overall superconducting gap projection, as illustrated in Figs. 3 (a), (b) and (c).
@@
$\Delta_{1/2}(k)=\sqrt{2(1+\lvert B_{\bm{k}}^{\pm}\lvert^{2})}\phi_{1k},$ (13)
with $B^{\pm}_{\bm{k}}=i(\sqrt{\lambda^{2}+h_{\bm{k}}^{2}\lambda_{R}^{2}}\pm\lambda_{R}h_{\bm{k}})/\lambda$ and $h_{\bm{k}}=\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$. In Fig. 3 (e) and (f), we plot $\Delta_{1/2}(k)$ along the $\Gamma-X$ and $\Gamma-M$ directions, respectively. The sign-change feature of the gap function at the two Fermi momenta is clearly visible. Such sign-change gap structure is consistent with the Fermi surface nesting and $\chi^{zz}(\bm{q})$ in Figs. 1 (c)-(e). Similar analysis can be performed for channel $B_{2}(d_{\pm})$, as shown in Figs. S8-S10[43]. The results indicate that its sign-change feature cannot come from the pure spin-singlet channel and depends on the mixing of spin-singlet and spin-triplet components[43], in contrast to the $A_{1}(S_{\pm})$ pairing.
@@
Topological superconductivity— For noncentrosymmetric systems with time-reversal symmetry, the criteria for identifying TSC require that a pair of spin-split Fermi surfaces enclose an odd number of time-reversal invariant momentum points, and the order parameter must have opposite signs on the two Fermi surfaces[45]. Additionally, the fully gapped superconducting state is preferable. According to this criterion, the Fermi surface in Fig. 1 (c) and the corresponding superconducting gap in Fig. 3 (a) or (b) support TSC.
@@
Figure 4: (a) The Wilson loop calculation for the $A_{1}(S_{\pm})$ superconducting state. (b) The quasi-particle spectrum of $A_{1}(S_{\pm})$ superconducting state. Note that the in-gap edge states are double-degeneracy with opposite chriality. Periodic and open boundary conditions are applied along $x$ and $y$ direction, respectively. In both (a) and (b), the parameters are the same as those in Fig. 3 (a).
@@
For the two-dimensional system with time-reversal symmetry considered here, the Chern number is always zero, and the topological nature of the superconducting state can be characterized by a topological invariant $Z_{2}$. Due to the lack of spatial inversion symmetry, the $Z_{2}$ invariant can be determined by calculating the Wannier centers of the superconducting state’s wave functions[43, 46] and verified through the computation of the edge spectrum[47, 48]. Figs. 4 (a) and (b) show the the Wannier center and edge spectrum, respectively, for the superconducting state with the Fermi surface from Fig. 1 (c) and the gap structure from Fig. 3 (b). It is straightforward to conclude that this superconducting state has a topological invariant $Z_{2}=1$. The superconducting state in $B_{2}(d_{\pm})$ channel has nodes, however, it is also topologically nontrivial[43, 49], and exhibites Majorana flat bands and dispersive Majorana states on edges.
@@
Outlook and conclusion— The TSC state driven by predominant $s$-wave spin-singlet pairing in our study has significant implications. First, TSC states realized through $p$- or $f$-wave spin-triplet pairing are extremely rare in real material systems. In contrast, almost all reported superconductors exhibit $s$-wave spin-singlet pairing, including cuprate, iron-based, and nickelate high-temperature superconductors[50, 51, 52]. Second, the splitting of Fermi surfaces accompanied by spin-texture characteristics due to broken inversion symmetry has been experimentally reported in systems such as cuprate and iron-based superconductors[30, 31]. For instance, spin splitting induced by Rashba SOC could potentially be tuned via electric fields or other external controls. Our proposal offers a viable pathway to realize bulk topological superconductivity in high-temperature superconducting systems without relying on proximity effects. If achieved, this could greatly advance the exploration of topological quantum computing based on topological superconductivity.
@@
In summary, we investigated superconductivity in a new class of correlated non-centrosymmetric systems. Using the theory of antiferromagnetic-fluctuation-mediated superconductivity, we identified four possible superconducting states $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ and $B_{2}(d_{\pm})$ under different model parameters. all of which are predominantly driven by spin-singlet pairing with $A_{1}$, $B_{1}$ and $B_{2}$ pairing symmetries. Notably, the superconducting state with $A_{1}(S_{\pm})$ IR is predominantly driven by spin-singlet pairing and belongs to a bulk topological superconducting state characterized by a $Z_{2}=1$ topological invariant. Our proposal is different from all present scenarios based on proximity effects and strong $p$- or $f$-wave spin-triplet pairing. We further highlighted the potential realization of such topological superconducting states in certain high-temperature superconductors, providing a new pathway for exploring next-generation high-temperature topological superconductors.

### v2

@@ IV Parity-mixed state and $S_{\pm}$-wave state
Once superconductivity is induced by antiferromagnetic fluctuations, the pairing symmetry of the superconducting state can be classified according to the IRs of the $C_{4v}$ group. Table 1 lists the possible pairings with their form factors up to the next-nearest neighbor and their corresponding IRs.
@@
By solving Eq. (10), the leading pairing can be determined. The phase diagrams concerning $\lambda_{R}-n$ with $n$ being the occupied particle number per site under various parameters are presented in Fig. 3. Across a wide-range parameter space, we identify two possible leading pairing channels: $A_{1}(S_{\pm})$, and $B_{2}$ as shown in Figs. 3 (a) and (b). Notably, the $A_{1}(S_{\pm})$ state exhibits a highly unconventional sign-change gap function, i.e., $S_{\pm}$-wave pairing, as shown in Figs. 2 (a), (f), (k), (p). The phase diagrams in Fig. 3 indicate that when the strength of RSOC is weak, the superconducting gap maintains a conventional $d$-wave gap, consistent with numerical results from the single-orbital model of cuprates. When the Hubbard $U$ is relatively small, a strong RSOC, leads to a inter-pocket sign-change superconducting gap as shown in Figs. 3 (a) and (b). When $n$ is close to 0.5, the $A_{1}$ IR with an $s$-wave inter-pocket sign-change superconducting gap is dominant, while for a little smaller $n$, the $d$-wave inter-pocket sign-change gap labeled by $d_{\pm}$ become prominent. Additionally, when $n$ is around 0.2, there exists a region of $A_{1}$ IR with spin-triplet dominant. In the case where Hubbard $U$ is large shown in Fig. 3 (c) and (d), spin density wave (SDW) states emerge in the large $n$ region, overshadowing the $A_{1}$ IR in large $n$ regime, resulting in a transition from spin-triplet-dominant to spin-singlet-dominant types in the $A_{1}$ gap function, as shown in Figs. 3 (c) and (d).
@@
Figure 4: (a)-(f) The plottings illustrate the contribution weights from six channels to the $A_{1}(S_{\pm})$-wave as functions as some specific parameters. The other fixed parameters are set as flollows: (a) $\lambda_{so}=3$, $n=0.26$ and $\lambda_{R}=0.35$, corresponding to Fig. 2 (p), (b) $\lambda_{so}=3$, $n=0.26$ and $U=7.5$, (c) $\lambda_{R}=0.35$, $n=0.26$ and $U=7.5$, (d) $\lambda_{so}=3$, $n=0.5$ and $\lambda_{R}=0.9$, (e) $\lambda_{so}=3$, $n=0.5$ and $U=3.5$, (f) $\lambda_{R}=0.725$, $n=0.5$ and $U=3.5$. Note that in (a), (b), (d) and (e), when the condition requiring the absolute value of the weight is not imposed, the pairing in $\tau^{3}\sigma^{0}$ channel can acquire a minus sign under a direct projection. In (f), the right part of the green-line minimum is also taken in absolute value for the same reason.
To investigate the parity-mixed states and elucidate the origin of this $S_{\pm}$-wave pairing, we express the gap function in a parameterized form as follows[46],
$\hat{\Delta}_{k}=i\sigma^{2}[\hat{\phi}_{k}+\bm{\hat{d}_{k}}\cdot\bm{\sigma}].$ (15)
In the one-orbital Rashba model, the projection of the gap function onto the two spin-split Fermi surfaces can be expressed in a simplified form, $\Delta_{1/2}(k)=\phi_{k}\pm|\hat{d}_{k}|$ with $\phi_{k}$ as a scalar. Thus, the aforementioned sign-change gap structure only occur when the triplet component is relatively strong and has helical $p$- or $f$-wave spin-triplet pairing, such as when the filling is near the van Hove singularity[32, 34, 35]. Any doping away from the van Hove singularity consistently results in predominant nodal d-wave spin-singlet pairing[33]. However, for the two-orbital model considered here, additional possibilities arise. We project the six sub-channels $i\sigma^{2}$$\Delta_{0}$($\tau^{0}\sigma^{0}\phi_{1k}$, $\tau^{3}\sigma^{0}\phi_{2k}$, $\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}$, $\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$, $\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}$, $\tau^{2}\sigma^{3}d_{6k}$) in the spin-orbital representation, listed in Table 1 onto the two Fermi surfaces. The details can be found in Appendix B. Note that the first two sub-channels belong to spin-singlet while the latter four correspond to spin-triplet.
We further distinguish $A_{1}(S_{\pm})$ states according to the weights of each channel with varying of some parameters, as shown in Fig. 4. This demonstrates that the superconducting state exhibits rich parity-mixed characteristics. Note that the emergence of $B_{2}(d_{\pm})$ state requires strong RSOC; an example with $\lambda_{R}=0.95$ is provided in Appendix C.
Figures 4 clearly demonstrate that the $S_{\pm}$-wave state exhibits parity-mixed, with the relative contributions of the each channel depending on specific model parameters. This tunability makes our model a promising platform for exploring parity-mixed superconductivity, should it be realized in future material systems. As shown in Figs. 4 (a) and (d), the ranking of the contributions from each channel is almost unaffected by Hubbard $U$; however, in certain cases where channel contributions are similar, Hubbard $U$ can alter the leading channel as shown in Fig. 4 (a). For nn=0.26, fig. 4 (b) illustrates that when $\lambda_{R}$ is small, the $S_{\pm}$-wave state is predominantly even-parity channels ($\tau^{0}\sigma^{0}$ and $\tau^{2}\sigma^{3}$). As $\lambda_{R}$ increases, their contribution diminishes, while the odd-parity components increase. This suggests that the spin-triplet pairing in this system is primarily induced by $\lambda_{R}$, fundamentally distinct from the intrinsic $p$-wave or $f$-wave triplet pairing mediated by ferromagnetic fluctuations near a van Hove singularity via doping. In contrast, our $S_{\pm}$-wave state is mediated by antiferromagnetic fluctuations, consistent with this mechanism. For nn=0.5 as shown in fig. 4 (e), the dominant behavior of even-parity channels becomes even more pronounced, and over a wide range of the parameter $\lambda_{R}$, the spin-singlet channel remains the leading. This further indicates that the origin of the $S_{\pm}$ state is unrelated to $p$-wave or $f$-wave pairing. Therefore, in both weak and strong RSOC regime, the gap function of the $S_{\pm}$-wave state can be mainly governed by the singlet component, regardless of the presence of triplet contributions—even though the triplet channels can produce the same gap sign distribution, as shown in Fig. A3 in Appendix. Figure 4 (c) and (f) indicates that $\tau^{0}\sigma^{0}$ can be enhanced by $\lambda_{so}$. As discussed in relation to Fig. 2, the introduction of $\lambda_{so}$ enhances the contribution of $\tau^{0}\sigma^{0}$ by canceling transverse susceptibility, ultimately strengthening inter-pocket interactions, which also enhances another even-parity channel, $\tau^{2}\sigma^{3}$.
V Topological superconductivity
@@
For noncentrosymmetric systems with time-reversal symmetry, the criteria for identifying TSC require that a pair of spin-split Fermi surfaces enclose an odd number of time-reversal invariant momentum points such as $\Gamma$, $X$ and $M$ points, as shown in Fig. 1 (b), where the Hamiltonian remains invariant under the time-reversal operation, and the order parameter must have opposite signs on the two Fermi surfaces[47]. Additionally, the fully gapped superconducting state is preferable. According to this criterion, the Fermi surface in Fig. 2 (p), with the $S_{\pm}$-wave pairing we focus on, support TSC.
@@
Figure 5: (a) The Wilson loop[48] calculation for the $S_{\pm}$ superconducting state. (b) The quasi-particle spectrum of $S_{\pm}$ superconducting state. Note that the in-gap edge states are double-degeneracy with opposite chriality. Periodic and open boundary conditions are applied along $x$ and $y$ direction, respectively. In both (a) and (b), the parameters are the same as those in Fig. 2 (p). The parameter $\Delta_{0}$ represents the coefficient of the form factors $\phi_{i,k}$ and $\bm{d_{j,k}}$ in the superconducting gap function and is set to 0.1.
@@
For the two-dimensional system with time-reversal symmetry considered here, the Chern number is always zero, and the topological nature of the superconducting state can be characterized by a topological invariant $Z_{2}$. Due to the lack of spatial inversion symmetry, the $Z_{2}$ invariant can be determined by calculating the Wannier centers of the superconducting state’s wave functions[48] and verified through the computation of the edge spectrum[49, 50]. The details can be found in Appendix B. Figs. 5 (a) and (b) show the Wannier center and edge spectrum, respectively, for the superconducting state with the Fermi surface from Fig. 2 (p) and the gap structure from Fig. A2. It is straightforward to conclude that this superconducting state has a topological invariant $Z_{2}=1$. The $B_{2}(d_{\pm})$ state has nodes, however, it is also topologically nontrivial[51], and exhibites Majorana flat bands and dispersive Majorana states on edges (See Appendix C).
VI Outlook and conclusion.
@@
The TSC state driven by predominant spin-singlet pairing in our study has significant implications. First, TSC states realized through $p$- or $f$-wave spin-triplet pairing are extremely rare in real material systems. In contrast, almost all reported superconductors exhibit spin-singlet pairing, including cuprate, iron-based, and nickelate high-temperature superconductors[52, 53, 54]. Second, the splitting of Fermi surfaces accompanied by spin-texture characteristics due to broken inversion symmetry has been experimentally reported in systems such as cuprate and iron-based superconductors[30, 31]. For instance, spin splitting induced by Rashba SOC could potentially be tuned via electric fields or other external controls. Our proposal offers a viable pathway to realize bulk topological superconductivity in high-temperature superconducting systems without relying on proximity effects. If achieved, this could greatly advance the exploration of topological quantum computing based on topological superconductivity.
@@
In summary, we investigated superconductivity in a new class of correlated non-centrosymmetric systems. Using the theory of antiferromagnetic-fluctuation-mediated superconductivity, we identified three possible superconducting states $A_{1}(S_{\pm})$, $B_{2}$ and $B_{2}(d_{\pm})$ under different model parameters. Notably, all states are parity-mixed, and the ratio between even-parity and odd-parity components can be tuned by the model parameters. More importantly, the $A_{1}(S_{\pm})$ state belongs to a bulk topological superconducting state characterized by a $Z_{2}=1$ topological invariant. In both weak and strong RSOC regime, the $A_{1}(S_{\pm})$ state can be predominantly driven by spin-singlet components, while $B_{2}(d_{\pm})$ is a nodal TSC and only emerges in strong RSOC regime. Our proposal is different from all present scenarios based on proximity effects and strong $p$- or $f$-wave spin-triplet pairing.